# Pressing Intensity: An Intuitive Measure for Pressing in Soccer

### v1

@@ Abstract
Pressing is a fundamental defensive strategy in football, characterized by applying pressure on the ball owning team to regain possession. Despite its significance, existing metrics for measuring pressing often lack precision or comprehensive consideration of positional data, player movement and speed. This research introduces an innovative framework for quantifying pressing intensity, leveraging advancements in positional tracking data and components from Spearman’s Pitch Control model. Our method integrates player velocities, movement directions, and reaction times to compute the time required for a defender to intercept an attacker or the ball. This time-to-intercept measure is then transformed into probabilistic values using a logistic function, enabling dynamic and intuitive analysis of pressing situations at the individual frame level. the model captures how every player’s movement influences pressure on the field, offering actionable insights for coaches, analysts, and decision-makers. By providing a robust and intepretable metric, our approach facilitates the identification of pressing strategies, advanced situational analyses, and the derivation of metrics, advancing the analytical capabilities for modern football.
@@ 2.1 Time to Intercept a Moving Object
The time to intercept an opponent or ball ($T_{i,j}$) estimates the time required for a defending player $i$ (the player applying pressure) to intercept an attacking target $j$ based on their respective locations ($\vec{r}_{i}(t)$ and $\vec{r}_{j}(t)$) and velocities ($\vec{v}_{i}(t)$ and $\vec{v}_{j}(t)$) at time $t$, a similar reaction time parameter, a similar maximum running velocity and an updated penalty parameter ($\tau_{\beta}$) as described in Formula 1).
@@
$\tau_{i,j}(t)=\frac{\|\vec{r}_{j}(t)+\vec{v}_{j}(t)-(\vec{r}_{i}(t)+\vec{v}_{i}(t)\cdot\tau_{r})\|}{v_{\text{max}}}$ (2)
$\tau_{\beta}(t)$ $=\frac{\|\vec{u}\|\cdot\beta}{\pi}$ (3)
where $\text{where }\quad\vec{u}$ $=(\vec{r}_{i}(t)+\vec{v}_{i}(t))-\vec{r}_{i}(t)$
$\vec{v}$ $=\vec{r}_{j}(t)+\vec{v}_{j}(t)$
$\beta$ $=\arccos\left(\frac{\vec{u}\cdot\vec{v}}{\|\vec{u}\|\|\vec{v}\|}\right)$

### v2

@@ Abstract
Pressing is a fundamental defensive strategy in football, characterized by applying pressure on the ball owning team to regain possession. Despite its significance, existing metrics for measuring pressing often lack precision or comprehensive consideration of positional data, player movement, and speed. This research introduces an innovative framework for quantifying pressing intensity, leveraging advancements in positional tracking data and components from Spearman’s Pitch Control model. Our method integrates player velocities, movement directions, and reaction times to compute the time required for a defender to intercept an attacker or the ball. This time-to-intercept measure is then transformed into probabilistic values using a logistic function, enabling dynamic and intuitive analysis of pressing situations at the individual frame level. The model captures how every player’s movement influences pressure on the field, offering actionable insights for coaches, analysts, and decision-makers. By providing a robust and interpretable metric, our approach facilitates the identification of pressing strategies, advanced situational analyses, and the derivation of derived metrics, advancing the analytical capabilities for modern football.
The full implementation of the the Pressing Intensity algorithm is available as open-source code through the unravelsports [14] Python package, introduced by Bekkers & Sahasrabudhe (2024) [3].
@@ 2.1 Time to Intercept a Moving Object
The time to intercept an opponent or ball ($T_{i,j}$) estimates the time required for a defending player $i$ (the player applying pressure) to intercept an attacking target $j$ based on their respective locations ($\vec{r}_{i}(t)$ and $\vec{r}_{j}(t)$) and velocities ($\vec{v}_{i}(t)$ and $\vec{v}_{j}(t)$) at time $t$, a similar reaction time parameter, a similar maximum running velocity, an updated penalty parameter ($\tau_{\beta}$) and a small positive value (e.g. $\epsilon=10^{-5}$) as described in Formula 1).
@@
$\tau_{i,j}(t)=\frac{\|\vec{d}\|}{v_{\text{max}}}\quad\text{where}\quad\begin{aligned} \vec{d}_{j}&=\vec{r}_{j}(t)+\vec{v}_{j}(t)\\ \vec{d}_{i}&=\vec{r}_{i}(t)+\vec{v}_{i}(t)\cdot\tau_{r}\\ \vec{d}&=\vec{d}_{j}-\vec{d_{i}}\end{aligned}$ (2)
$\tau_{\beta}(t)=\frac{\|\vec{u}\|\cdot\beta}{\pi}\quad\text{where }\quad\begin{aligned} \vec{u}&=(\vec{r}_{i}(t)+\vec{v}_{i}(t))-\vec{r}_{i}(t)\\ \vec{v}&=\vec{d}_{j}-\vec{r}_{i}(t)\\ \beta&=\arccos\left(\frac{\vec{u}\cdot\vec{v}}{\|\vec{u}\|\cdot\|\vec{v}\|+\varepsilon}\right)\end{aligned}$ (3)